A quiz with 4 multiple choice questions, each with 3 choices, was given. A student needs either 3 or 4 right to pass the quiz, but is totally clueless like myself. What is the probability of passing by pure guesswork? I wonder if my teacher made this problem up to make fun of me? :(.. thankyou Shelly
Answer by scott8148(6628)

there is only one way to get 4 right, and the probability is (1/3)^4

there are 4 ways to get 3 right (each of 4 questions being the only one wrong)
__ the probability is 4[((1/3)^3)(2/3)]

(1/3)^4+4[((1/3)^3)(2/3)]=1/81+4(2/81)=1/81+8/81=9/81=1/9

one chance in nine of guessing your way through the test